# MapFormer: Self-Supervised Learning of Cognitive Maps with Input-Dependent Positional Embeddings

### v1

@@ 1 Introduction
Figure 1: Cognitive Maps: Disentangling Structure and Content for Path Integration in Episodic and Working-Memory Models. (a) Cognitive maps model the relationships between entities, such as objects in space or people working in an organization. In a cognitive map, both the entities and their position are represented but they can be factorized from each other. The position (or, more generally, the ‘state’) in a cognitive map is updated via actions (blue arrows). For example, actions can update position on a grid (e.g., ‘step right’ in the left drawings), or in more abstract environments, on an organization chart (e.g. ‘the boss of’, on the right). Actions are assumed to have the same effect on the state in the map across all positions and across all environments (up vs bottom). Path integration can be achieved by sequentially combining actions together $a_{0\to T}=(a_{0}\to...\to a_{T})$ (green arrow). (b) (Top) The mental trajectory of a rodent navigating a maze lies on a torus (figure adapted from Gardner2022 under $C$ BY 4.0 license); (bottom) trajectories during a 2D navigation task in MapFormers also lie on a torus. The red curve represents a square trajectory performed by the simulated agent during a navigation task, returning to its initial starting point, as desired. (c-d) An illustration of the conceptual framework behind MapFormers, and of the differences between the two variants for Episodic Memory (EM) and Working Memory (WM) models Whittington2025. (c) Episodic Memory (EM) model: Positions (or ‘states’, more generally) in a cognitive map $p_{t}$’s (in blue), at times $t$’s, are encoded in a separate neural space by, e.g., a dedicated neural population. Such population does not directly interact with the input $x_{t}$ (red), but its state is updated via actions only. Actions $a_{t}$’s, when they occur in the input, can update the position via input-dependent matrices $W_{a_{t}}$ (green): $p_{t}=W_{a_{t}}p_{t-1}$. $W_{a}$’s are input-dependent matrices and are computed through an exponential map, which maps between abstract representations of actions $A$’s (residing in the Lie-Algebra space) and the corresponding weight matrix $W_{a}$’s (in the corresponding Lie group; section 3.4). The exponential mapping is computed based on the integration time $\Delta_{t}$ of the action (light green) – roughly, for how long the action needs to be applied. Finally, the conjunction (purple) of the content (the ‘what’) and the structure (the ‘where’) can be computed. (d) Working Memory (WM) model: In contrast to the EM model, which factorizes content (red) and position (blue) in two separate neural spaces, MapFormers-WM represent position implicitly. Actions in the WM model directly update the conjunction of what and where (purple). The conjunction states $h_{t}$’s in the WM model are thus smaller by a quadratic factor compared to the EM model.
@@
The ability of Transformers NIPS2017_3f5ee243 to process sequences in parallel has enabled the training of Large Language Models (LLMs) on increasingly large datasets, creating models able to write poetry, solve complex mathematical problems, or even describe what they observe in an image, results so impressive that debates about their reaching AGI (Artificial General Intelligence) have already started bubeck2023sparksartificialgeneralintelligence; morris2024levelsagioperationalizingprogress. Yet, the same models learn incoherent world models vafa2024evaluatingworldmodelimplicit, make logical mistakes on tasks that are trivial to us humans williams2024easyproblemsllmswrong, and even when they are right, they can be so for the wrong reasons mccoy-etal-2019-right. Moreover, the fact that different architectures lie at different levels within Chomsky’s hierarchy of language deletang2023neural suggests that a key structural bias is still missing in order to reach human-like generalization abilities and AGI.
@@
In comparison, human and animal cognitive abilities appear much more robust and flexible. It has been suggested that the reason for this is our ability to build cognitive maps BEHRENS2018490, internal models of the relationships among entities or events in the world (Figure 1a). This definition was first introduced in 1948 by Edward Tolman Tolman1948-TOLCMI as a way to explain how rats could successfully find new routes in a maze after old ones were blocked, behavior that cannot be explained by simple conditioning, suggesting that rats develop an internal cognitive map of their environment to flexibly adapt to new scenarios. Neural evidence of internal maps was first discovered in the hippocampus O'Keefe1978, with place cells firing every time a rodent came back to a specific location. Later, grid cells were found in the entorhinal cortex (EC) Hafting2005, firing periodically with a “grid-like” pattern as an animal was exploring a room, adding further evidence of an internal map within mammal’s brains.
@@
Cognitive maps are essential for generalization and flexible decision-making, as they provide an internal model that represents the structural relationships among entities in the world. This abstraction holds true for both spatial navigation (e.g., finding a shortcut to get back home, or finding a detour around an obstacle) and non-spatial, conceptual information (e.g., inferring family relations). For example, in the case of spatial navigation, if a wall blocks the road ahead, understanding that the sequences $(\rightarrow,\rightarrow)$ and $(\uparrow,\rightarrow,\rightarrow,\downarrow)$ are equivalent allows an agent to ignore unexpected obstacles by simply walking around before continuing on the desired path. If someone says “Bob is the brother of Alice’s mother,” when asked “Who is Alice’s uncle?” one can easily reply “Bob” by simply knowing the abstract relationships of a family tree. Seeing structure as a connected graph of relations unifies spatial and non-spatial cognitive maps, and is the reason why some have theorized that the neural mechanism allowing the creation of spatial cognitive maps might be “recycled” in humans to build abstract knowledge graphs Aronov2017; KRIEGESKORTE2016280; Park2021 (fig. 1(a)).
@@
Cognitive maps achieve robust generalization by separating the abstract, invariant structure of the world from the transient, variable content of observations. Because cognitive maps define the position of an object via its relations to other entities in the world, these models can be formalized as graphs, whose nodes represent objects and edges the relations connecting them Whittington770495. This factorizes what has been identified from where it was seen. Making structure independent from content dramatically helps generalization, assuming that the structural relations learned in one environment remain true in new ones. Indeed, making a step right then left brings us to the same point whether we perform it in Paris or New-York, while our representation of numbers remains the same whether we are counting apples or oranges. Disentangling structure from content also decreases the required amount of training data by a quadratic factor, since the joint distribution can be expressed as the product of its two marginals: $P(x_{i},p_{j})=P_{x}(x_{i})P_{p}(p_{j})$. In a world where structure remains stable but observations are infinite, this factorization is fundamental.
@@
What kind of structural bias must be introduced into a deep learning architecture to enable it to effectively learn a cognitive map and perform robust path integration? Previous work made progress in modeling cognitive maps with RNNs with action-dependent matrices Whittington770495, allowing them to solve structure-dependent tasks and perfectly generalize to unseen scenarios. This work also explains the duality between episodic and working memory (EM / WM) Whittington2025: EM uses action-dependent matrices to sequentially update a separate pool of neurons representing the agent’s abstract location in a cognitive map, used for long term recall of structured memories stored in synaptic weights (fig.1(c)); WM uses these matrices to directly move the state into distinct neural subspaces — acting as memory slots — allowing the model to keep the whole structure in neural activity (fig.1(d)). Although very promising, these models do not scale since (1) their action-dependent matrices need to be sequentially multiplied to perform path integration, making them slow to train and incompatible with parallel sequence processing, and (2) the models can only be applied to problems where the underlying structure is already known (for example, where the number of actions is known rather than learned).
To address this, we introduce MapFormers, a class of transformers that learn context-dependent cognitive maps with a generic neural mechanism, allowing the learning of structure at larger scale. MapFormers compute path integration in parallel and therefore allows scaling up to complex problems. Furthermore, MapFormers are trained in a fully self-supervised manner, and learn by themselves to disambiguate between two types of tokens (actions and observations), which have opposite effects on the learned representations. The former update the internal state’s structure via learned operations but leave its semantic content untouched while the latter leave the internal structure fixed but change the state’s content. Empirically, we argue that MapFormers, to the best of our knowledge, are the first generic architectures to learn cognitive maps and generalize to longer sequences without supervision, and with minimal assumptions about the underlying structure. We introduce two variants of MapFormers, which unify relative rotary (RoPE su2023roformerenhancedtransformerrotary) and absolute positional embeddings, respectively as models of working and episodic memory. We confirm the fundamental role that action-dependent positional embeddings play in encoding the location of tokens within a context-dependent cognitive map whittington2022relatingtransformersmodelsneural. Finally, we draw theoretical arguments from Lie-group theory to characterize the class of structures that sequence-processing models are able to learn, allowing us to show that diagonal matrices used in current selective State Space Models are not expressive enough to learn cognitive maps.
@@ 2 Related Literature
In seminal work Whittington770495, the authors presented a set of tasks, for both spatial (navigation) and non-spatial (e.g., family trees) problems, which require learning a cognitive map in order to achieve strong generalization abilities. The Tolman-Eichenbaum Machine (TEM) was suggested as a model that can learn cognitive maps to solve these tasks. Specifically, TEM is an RNN-based architecture that sequentially updates its internal position with action-dependent weights, before querying a Hopfield memory module ramsauer2021hopfieldnetworksneed with its updated position to retrieve what has been seen at a previously visited location.
@@
However, a key limitation of models based on standard RNNs — including both the EM and WM architectures — for large-scale applications is the inherent computational bottleneck created by their sequential processing. This is particularly restrictive in real-world scenarios, such as long-range path integration, where the cumulative effect of a sequence of actions requires the slow, iterative multiplication of recurrent weight matrices. It is best illustrated in TEM-t whittington2022relatingtransformersmodelsneural, where it is shown that TEM can be seen as a Transformer, which however loses its parallel sequence processing abilities due to the sequential update of positional embeddings.
To keep path integration compatible with the parallel processing ability of Transformers, we take inspiration from the input dependent matrices of modern selective State Space Models gu2024mambalineartimesequencemodeling and their established links with Transformers dao2024transformersssmsgeneralizedmodels to enable path integration in parallel and update all positional embeddings at once. We also introduce theoretical arguments from Lie-group theory to characterize the class structure that a model can learn based on the characteristics of its recurrent matrix, showing that the diagonal matrices used in Mamba models — necessary to scale these architectures — are not expressive enough to learn cognitive maps. From this, we propose two new architectures based on the Transformer framework for learning both EM-based and WM-based cognitive maps in a fully self-supervised setting (fig. 1(c) and 1(d), fig. 2).
@@
Figure 2: MapFormers: Path-Integration Transformer-Based Models that Learn Cognitive Maps. Overview of our Transformer architectures unifying working and episodic memory as relative and absolute positional encodings, respectively. In both case, rotation angle $\theta$ is obtained via a low-rank projection of input $X$, before applying a cumulative sum (cumsum) along the temporal dimension, to perform path integration (PI) in parallel $\theta^{PI}=\omega\Delta^{PI}=\omega\sum_{s\leq T}\Delta_{s}$. Input-dependent block-diagonal matrices computed via $R_{\theta^{PI}}=\exp({\theta^{PI}}S),\,S:=-S^{\top}$ capture the task’s latent structure, while diagonal matrices used in selective SSMs do not. (a) Working Memory MapFormer (MapWM) with relative positional encoding, where $Q,K$ are rotated via input-dependent matrix $R_{\theta_{PI}}$. In MapFormers, $\theta$ is learned, but keeping it fixed gives exactly RoPE (b) Episodic Memory MapFormer (MapEM) with two parallel attentions $A_{X}$ and $A_{P}$. The former captures raw similarities between external inputs and the latter encodes the problem’s structure. Combining these two attention matrices with element-wise multiplication $A:=A_{X}\odot A_{P}$ is equivalent to computing a single attention $A_{G}$ where keys and queries are the conjunction of observation and position: $g_{t}=\text{vec}(x_{t}^{\top}.p_{t})\in\mathbb{R}^{d^{2}}$, like in TEM-t whittington2022relatingtransformersmodelsneural. In MapEM, absolute position embeddings is obtained via path integration of initial and shared key/query positions $Q_{P}^{*}$ and $K_{P}^{*}$.
@@ 3 General Setup
This section provides the theoretical details that motivate the development of MapFormers, as well as their implementation details. Section 3.1 describes the tasks we used for evaluating the models: (1) A selective-copy task and (2) Forced-Navigation tasks. Section 3.3 describes the framework for action-dependent matrices and their formulation as a group. Section 3.4 introduces the idea that for learning a cognitive map, path integeration can be more efficiently computed at the Lie algebra vector space, rather than at the group level. Finally, given these insights from Lie theory, 3.5&3.6 show how to build state-space models (SSMs) and Transformer-based models, respectively, that can learn a cognitive map.

### v2

@@ 1 Introduction
Figure 1: Cognitive Maps: Disentangling Structure and Content for Path Integration in Episodic and Working-Memory Models. (a) Cognitive maps model the relationships between entities, such as objects in space or people working in an organization. In a cognitive map, both the entities and their position are represented but they can be factorized from each other. The position (or, more generally, the ‘state’) in a cognitive map is updated via actions (blue arrows). For example, actions can update position on a grid (e.g., ‘step right’ in the left drawings), or in more abstract environments, on an organization chart (e.g. ‘the boss of’, on the right). Actions are assumed to have the same effect on the state in the map across all positions and across all environments (up vs bottom). Path integration can be achieved by sequentially combining actions together $a_{0\to T}=(a_{0}\to...\to a_{T})$ (green arrow). (b) (Top) The mental trajectory of a rodent navigating a maze lies on a torus (figure adapted from Gardner2022 under $C$ BY 4.0 license); (bottom) trajectories during a 2D navigation task in MapFormers also lie on a torus. The red curve represents a square trajectory performed by the simulated agent during a navigation task, returning to its initial starting point, as desired. (c-d) An illustration of the conceptual framework behind MapFormers, and of the differences between Episodic Memory (EM) and Working Memory (WM) models Whittington2025. (c) Episodic Memory (EM) model: Positions (or ‘states’, more generally) in a cognitive map $p_{t}$’s (in blue), at times $t$’s, are encoded in a separate neural space by a dedicated neural population, which does not directly interact with the input $x_{t}$ (red), being updated via actions only. Actions $a_{t}$’s, when they occur in the input, can update the position via input-dependent matrices $W_{a_{t}}$ (green): $p_{t}=W_{a_{t}}p_{t-1}$. $W_{a}$’s are input-dependent matrices and are computed through an exponential map, which maps between abstract representations of actions $A_{t}:=\Delta_{t}\omega S=\theta_{t}S$ (in the Lie-Algebra) and the corresponding weight matrix $W_{a_{t}}:=\exp(A_{t})=R_{\theta_{t}}$ (in the corresponding Lie group; section 3.3). The exponential mapping is computed based on the integration time $\Delta_{t}$ of the action (light green) – roughly, how long the synaptic weights $S$ need to be applied. Finally, the conjunction (purple) of the content (what) and the structure (where) can be computed. (d) Working Memory (WM) model: In contrast to the EM model, which factorizes content (red) and position (blue) in two separate neural spaces, MapFormers-WM represent position implicitly. Actions in the WM model directly update the conjunction of what and where (purple). The conjunction states $h_{t}$’s in the WM model are thus smaller by a quadratic factor compared to the EM model.
@@
The ability of Transformers NIPS2017_3f5ee243 to process sequences in parallel has enabled the training of Large Language Models (LLMs) on increasingly large datasets, creating models able to write poetry, solve complex mathematical problems, or even describe what they observe in an image, results so impressive that debates about their reaching AGI (Artificial General Intelligence) have already started bubeck2023sparksartificialgeneralintelligence; morris2024levelsagioperationalizingprogress. Yet, the same models learn incoherent world models vafa2024evaluatingworldmodelimplicit, make logical mistakes on tasks that are trivial to us humans williams2024easyproblemsllmswrong, and even when they are right, they can be so for the wrong reasons mccoy-etal-2019-right. Moreover, the fact that Transformers fail to generalize OOD to non-regular formal tasks deletang2023neural shows that some key structural biases are most likely missing for transformer-based models to reach human-like generalization abilities and AGI.
@@
In comparison, human and animal cognitive abilities appear much more robust and flexible. It has been suggested that the reason for this is our ability to build cognitive maps BEHRENS2018490, internal models of the relationships among entities or events in the world (Figure 1a). This notion was first introduced in 1948 by Edward Tolman Tolman1948-TOLCMI as a way to explain how rats could successfully find new routes in a maze after old ones were blocked, behavior that cannot be explained by simple conditioning, suggesting that rats develop an internal cognitive map of their environment to flexibly adapt to new scenarios. Neural evidence of internal maps was first discovered in the hippocampus O'Keefe1978, with place cells firing every time a rodent came back to a specific location. Later, grid cells were found in the entorhinal cortex (EC) Hafting2005, firing periodically with a “grid-like” pattern as an animal was exploring a room, adding further evidence of an internal map within mammal’s brains.
@@
Cognitive maps could be important for generalization and flexible decision-making, as they provide an internal model that represents the structural relationships among entities in the world. This abstraction holds true for both spatial navigation (e.g., finding a shortcut to get back home, or finding a detour around an obstacle) and non-spatial, conceptual information (e.g., inferring family relations). For example, in the case of spatial navigation, if a wall blocks the road ahead, understanding that the sequences $(\rightarrow,\rightarrow)$ and $(\uparrow,\rightarrow,\rightarrow,\downarrow)$ are equivalent allows an agent to ignore unexpected obstacles by simply walking around before continuing on the desired path. If someone says “Bob is the brother of Alice’s mother,” when asked “Who is Alice’s uncle?” one can easily reply “Bob” by simply knowing the abstract relationships of a family tree. Seeing structure as a connected graph of relations unifies spatial and non-spatial cognitive maps, and is the reason why some have theorized that the neural mechanism allowing the creation of spatial cognitive maps might be “recycled” in humans to build abstract knowledge graphs Aronov2017; KRIEGESKORTE2016280; Park2021 (fig. 1(a)).
@@
Cognitive maps achieve robust generalization by separating the abstract, invariant structure of the world from the transient, variable content of observations BEHRENS2018490. Because cognitive maps define the position of an object via its relations to other entities in the world, these models can be formalized as graphs, whose nodes represent objects and edges the relations connecting them Whittington770495. This factorizes what has been identified from where it was seen. Making structure independent from content dramatically helps generalization, assuming that the structural relations learned in one environment remain true in new ones. Indeed, making a step right then left brings us to the same point whether we perform it in Paris or New-York, while our representation of numbers remains the same whether we are counting apples or oranges. Disentangling structure from content also decreases the required amount of training data by a quadratic factor, since the joint distribution can be expressed as the product of its two marginals: $P(x_{i},p_{j})=P_{x}(x_{i})P_{p}(p_{j})$. In a world where structure remains stable but observations are infinite, this factorization is fundamental.
@@
What kind of structural bias must be introduced into a deep learning architecture to enable it to effectively learn a cognitive map and perform robust path integration? Previous work made progress in modeling cognitive maps with RNNs with action-dependent matrices Whittington770495, allowing them to solve structure-dependent tasks and generalize to unseen scenarios. This work also explains the duality between episodic and working memory (EM / WM) Whittington2025: EM uses action-dependent matrices to sequentially update a separate pool of neurons representing the agent’s abstract location in a cognitive map, used for long term recall of structured memories stored in synaptic weights (fig.1(c)); WM uses these matrices to directly move the state into distinct neural subspaces — acting as memory slots — allowing the model to keep the whole structure in neural activity (fig.1(d)). Although very promising, these models do not scale since (1) their action-dependent matrices need to be sequentially multiplied to perform path integration, making them slow to train and incompatible with parallel sequence processing, and (2) the models can only be applied to problems where the underlying structure is already known (for example, where the number of actions is known rather than learned).
To address this, we introduce MapFormers, a class of transformers that learn context-dependent cognitive maps with a generic neural mechanism, allowing the learning of structure at larger scale. MapFormers compute path integration in parallel and therefore allows scaling up to complex problems. Furthermore, MapFormers are trained in a fully self-supervised manner, and learn by themselves to disambiguate between actions and observations, which have opposite effects on the learned representations. The former update the internal state’s structure via learned operations but leave its semantic content untouched, while the latter leave the internal structure fixed but change the state’s content. Empirically, we argue that MapFormers, to the best of our knowledge, are the first generic architectures to learn cognitive maps and generalize to longer sequences without supervision, and with minimal assumptions about the underlying structure. We introduce two variants of MapFormers, which unify relative rotary (RoPE su2023roformerenhancedtransformerrotary) and absolute positional embeddings, respectively as models of working and episodic memory, confirming the fundamental role that action-dependent positional embeddings play in encoding the location of tokens within a context-dependent cognitive map whittington2022relatingtransformersmodelsneural. Finally, we draw theoretical arguments from Lie-group theory to characterize the class of structures that sequence-processing models are able to learn, allowing us to show that diagonal matrices used in current selective State Space Models are not expressive enough to learn cognitive maps.
@@ 2 Related Literature
In a seminal work Whittington770495, the authors presented a set of tasks, for both spatial (navigation) and non-spatial (e.g., family trees) problems, which require learning a cognitive map in order to achieve strong generalization abilities. The Tolman-Eichenbaum Machine (TEM) was suggested as a model that can learn cognitive maps to solve these tasks. Specifically, TEM is an RNN-based architecture that sequentially updates its internal position with action-dependent weights, before querying a Hopfield memory module ramsauer2021hopfieldnetworksneed with its updated position to retrieve what has been seen at a previously visited location.
@@
However, a key limitation of models based on standard RNNs — including both the EM and WM architectures — for large-scale applications, is the inherent computational bottleneck created by their sequential processing. This is particularly restrictive in real-world scenarios, such as long-range path integration, where the cumulative effect of a sequence of actions requires the slow, iterative multiplication of recurrent weight matrices. This is true also for more recent work using Transformers - In a follow-up work, whittington2022relatingtransformersmodelsneural introduced TEM-t, a model for TEM that can be seen as a Transformer, where positional embeddings in the model represent abstract positions in a cognitive map. However, TEM-t cannot leverage parallel sequence processing abilities of Transformers since it requires sequential updates of positional embeddings with input-dependent matrices. MapFormers overcome this limitation by employing cumulative summation across the entire sequence, allowing the computation of path integration to occur in parallel, as we will see below.
Positional embeddings were traditionally added to Transformers because the attention mechanism is permutation invariant, with position either added explicitly within a separate pool of neurons NIPS2017_3f5ee243, or implicitly by representing the relative distance between keys and queries su2023roformerenhancedtransformerrotary; shaw2018selfattentionrelativepositionrepresentations. Positional embeddings factorize content from position in transformers, but since these positional representations are shared by all sequences, and are not input-dependent, they cannot represent the task’s latent structure. In line with TEM-t, other work has demonstrated the generalization power of sequence-dependent positional encoding. Specifically, Contextualized equivariant Position Encoding (TAPE) zhu2025rethinkingaddressinglanguagemodels introduced input-dependent MLP after the attention mechanism, updating positional embeddings for the next layer, which reverses the logic of TEM-t, where positions are modified before querying the attention module. CoPE golovneva2024contextualpositionencodinglearning modifies its positional encoding by compressing the relative distances between keys and queries, achieving almost perfect length generalization in some 1-dimensional tasks. PoPE gopalakrishnan2025decouplingwhatwherepolar follows a logic similar to CoPE, by directly modifying RoPE’s relative distances, meaning that in both CoPE and PoPE, relative distances are not constrained to be integers and can take any positive real value, allowing the model to shorten the distance between structurally similar tokens. While these three methods (TAPE, CoPE and PoPE) yield superior results in length generalization and a selection of formal tasks, their representation power is poorly understood. Specifically, little is known about what kind of structural features these model can or cannot learn. Also, no connections have been made to cognitive maps and path integration, which is one of the main goals of MapFormers.
@@
Figure 2: MapFormers: Path-Integration Transformer-Based Models that Learn Cognitive Maps. Overview of our Transformer architectures unifying working and episodic memory as relative and absolute positional encodings, respectively. In both case, rotation angle $\theta$ is obtained via a low-rank projection of input $X$, before applying a cumulative sum (cumsum) along the temporal dimension, to perform path integration (PI) in parallel $\theta^{PI}=\omega\Delta^{PI}=\omega\sum_{s\leq T}\Delta_{s}$. Input-dependent block-diagonal matrices computed via $R_{\theta^{PI}}=\exp({\theta^{PI}}S)$ ($S:=-S^{\top}$) capture the task’s latent structure, while diagonal matrices used in selective SSMs do not. (a) Working Memory MapFormer (MapWM) with relative positional encoding, where $Q,K$ are rotated via input-dependent matrix $R_{\theta_{PI}}$. In MapFormers, $\theta$ is learned, but keeping it fixed in MapWM gives exactly RoPE (b) Episodic Memory MapFormer (MapEM) with two parallel attentions $A_{X}$ and $A_{P}$. The former captures raw similarities between external inputs and the latter encodes the problem’s structure. Combining these two attention matrices with element-wise multiplication $A:=A_{X}\odot A_{P}$ is equivalent to computing a single attention $A_{G}$ where keys and queries are the conjunction of observation and position: $g_{t}=\text{vec}(x_{t}^{\top}.p_{t})\in\mathbb{R}^{d^{2}}$, like in TEM-t whittington2022relatingtransformersmodelsneural. In MapEM, absolute position embeddings is obtained via path integration of initial and shared key/query positions $Q_{P}^{*}$ and $K_{P}^{*}$.
@@ 3 General Setup
This section provides the theoretical details that motivate the development of MapFormers, as well as their implementation details. Section 3.1 describes the tasks we used for evaluating the models: (1) A selective-copy task and (2) Forced-Navigation tasks. Section 3.2 describes the framework for action-dependent matrices and their formulation as a group. Section 3.3 introduces the idea that for learning a cognitive map, path integeration can be more efficiently computed at the Lie algebra vector space, rather than at the group level. Finally, given these insights from Lie theory, 3.4&3.5 show how to build state-space models (SSMs) and Transformer-based models, respectively, that can learn a cognitive map.